You have exactly $537 to spend on party gifts for your rich uncle’s birthday party. You decide to get watches for the ladies (at $27.98 each), and beepers for the men (at $23.46 each). You know that the number of watches required will be 3 times as much as the number of beepers. How many of each item do you buy?
Answer by Mathtut(3670)

27.98W+23.46B=537 where W and B is number of watches and beepers respectively
W=3B
so 27.98(3B)+23.46B=537
83.94B+23.46B=537
107.4B=537
B=5
so W=3(5)=15